Chair: Ahmad

Next meeting chair: Simin

Attended: Lim, David Schunck, Jamie, Eloise.

Apologies: Simin, David Smith, Prad, Earl, Mars

General Issues:

Ke receiver status: Back to work

IFF sampler (DBBC3Hb) status: Nothing changed, the physically is fine, cooling is fine. it is a severe problem for observing. We have a potential workaround, switch the connections. A proper diagnosis from Bonn is needed.

Handover issues:

r41069

Hobart 12m

0024 UT: IFF (Core3h[6 has gone bad again. (JMc) 0028 UT sChedule halted to investigate/repair (JMc) 0148 UT Schedule resumed. Issue potentially identified as a loose connection to the ADB board. Configured ok now. (JMc) 0158 UT First good data (JMc) 1327UT Sampler F gone bad again, power cycled dbbc3 but still gave bad readings. Left until end of experiment for Jamie to tend to in the morning (no good readings after 1327UT) (Eloise)

crd121 Yarragadee 12m

I still cannot connect to ops8 (Jasper)

Experiment start delayed due to pcfs issues (fs was having some kind of connection issue with “lost connection” message printed in fs output). Attempting to restart the fs seemed to raise another issue where all the terminals froze and new ones couldn't be opened - it would spawn and then just disappear (Jasper). Eventually pcfsyg rebooted but dbbc then wouldn't connect properly so rebooted that too (Jasper). 1920UT experiment finally started (Jasper).

aua092

Could not connect to ops8 machine. Turned off/broken on-site?

Hobart 12m

1730UT Experiment started ok (Prad) The sampler value still bad. IFF delay bad, suspect hardware (w)

Katherine 12m

1812UT A flurry of 'antenna stuck'. Reset the drives and ran the 'operate option'. It was fixed by 1825UT. Halted the schedule for 2-3 minutes in this interim 1822UT-1824UT. Lost data between 1820-1830UT. (Prad). Occasional tracking errors show up (>0.045). Something to look into?2054

Yarragadee 12m

2030-2040 UT mk5 crash on module change (warren) 2054 UT resumed schedule (w)

R11069

Hobart 12m

1700UT Experiment not started at Hb due to DBBC fault. JMc was actively working on remedying the fault at the experiment start time (Mars) 1714UT Experiment began with a Sampler (#6 X-band, Y-pol) out of order, awaiting correction from on-site.

AUM054

Hobart 12m

started getting delay difference errors. difference is around -0.7us atm, so better not to restart dbbc again given that that doesn’t solved the issue last time. Note to observers: keep an eye for delay diff getting larger. Given a call to Jamie/ restart dbbb if goes beyond 1us (Jay).

Check the values are updating. Restating ups1hb.

AUM053

Hobart 12m

1926 UT Experiment started late without S-band, cannot access pduhb (Lim)

Open it from ops8 chromium

1940 UT pduhb turn out to be a browser issue. Now with S-band OK, technically missed first 10 scans (Lim). 1719 UT observed dbbc3hb f sampler going from “bad” to “good” to “bad” again. I am going to restart dbbc3hb. Last scan before powercycling dbbc3hb: 260-1719 (1057-797)(David) 1843 UT First scan after restart: 260-1843 (2052-474). I kept the dbbc off for about 30 minutes. f sampler is happy again (David)

t2155

Yarragadee 12m

1758 UT Experiment started late (missed first 5 scans) due to some issues with mk5 - but disk swap, on-site reboot (spoke w Martin there) and a few re-syncs w ddbc got it sorted. (Earl).